# Warm Inflation with Barrow Holographic Dark Energy

### v1

@@ 3.2 Model comparison with observational data
Figure 8: The figure illustrate the behaviour of $\frac{T}{H}$ with respect to $N$. The parameters chosen here are $\sigma_{r}=1$, $C_{1}=30^{5}$, $\gamma=5.04$, $\alpha_{0}=0.001$, $\beta_{0}=20^{-3}$, $m=2$.
@@ 4 Discussion and Conclusion
Warm inflation offers a framework for comprehending the dynamics of the early universe in contrast to classical cold inflation. The inflationary paradigm is introduced to rectify the shortcomings of the traditional cosmological model. Inflationary scenarios can be classified into two categories: warm inflation and cold inflation. In cold inflation, the matter field does not interact with radiation and it slowly rolls to its flat potential. However, for the warm inflationary scenario, the inflaton interacts with other fields resulting in the transmission of energy from the inflaton to the radiation field during slow-roll. The inflaton completely decays into radiation when inflation comes to an end, preventing inflation from causing the universe to enter a very cold phase. As a result, the universe enters into radiation radiation-dominated phase without the need for a separate reheating phase.
In this article, we have studied an inflationary scenario assuming that a holographic dark fluid is the source of inflation. We have chosen Barrow holographic dark energy for this purpose in a scenario where holographic dark energy interacts with radiation and energy transmits from holographic dark energy to radiation. Here we have revised the inflationary scenario in the high dissipative regime ($Q>1$). Assuming this condition we have reconstructed the Hubble parameter as a function of the e-folding number $N$. Slow roll conditions play an important role in warm inflation. A collection of slow-roll parameters determines how consistent the slow-roll approximation is. We have shown that these parameters will satisfy the warm inflationary conditions to validate our model. Moreover, after checking the tendencies of the different inflationary parameters, it was found that there is a good agreement with the observational data [53].
@@ Acknowledgments
P.R. acknowledges the Inter-University Centre for Astronomy and Astrophysics (IUCAA), Pune, India for granting visiting associateship.

### v2

@@ 3.2 Results
Figure 9: The figure illustrate the behaviour of $\frac{T}{H}$ with respect to $N$. The parameters chosen here are $\sigma_{r}=1$, $C_{1}=1.2$, $\gamma=1.02$, $\alpha_{0}=0.06$, $\beta_{0}=0.01$
Figure 10: The figure illustrate the behaviour of $Q$ with respect to $N$. The parameters chosen here are $\sigma_{r}=1$, $C_{1}=1.2$, $\gamma=1.02$, $\alpha_{0}=0.06$, $\beta_{0}=0.01$, $m=1$.
In Fig.(10) we have plotted dissipative parameter $Q$ with respect to $N$ and we can conclude that $Q>1$ is perfectly satisfied. This indicates a strong dissipative regime.
Table 4: The table shows numerical values of $n_{s}$, $r$ for different values of the parameters.
  $N$   $\Delta$   $C_{1}$   $\gamma$   $c_{t}$   $\alpha_{0}$   $\beta_{0}$   $n_{s}$   $r$
65 0.044 2.2 0.0004 0.5 0.02 0.002 0.9591 $2.115\times 10^{-99}$
65 0.046 2.2 0.0004 0.5 0.02 0.002 0.9675 $1.667\times 10^{-97}$
65 0.048 2.2 0.0004 0.5 0.02 0.002 0.9747 $1.441\times 10^{-95}$
65 0.065 2.2 0.0004 0.5 0.02 0.002 1.002 $1.162\times 10^{-82}$
65 0.070 2.2 0.0004 0.5 0.02 0.002 1.004 $3.289\times 10^{-79}$
65 0.075 2.2 0.0004 0.5 0.02 0.002 1.006 $3.563\times 10^{-76}$
60 0.048 2.1 0.0005 0.5 0.02 0.002 0.9445 $2.702\times 10^{-97}$
60 0.044 2.1 0.0005 0.5 0.02 0.002 0.9538 $9.232\times 10^{-96}$
60 0.046 2.1 0.0005 0.5 0.02 0.002 0.9623 $3.814\times 10^{-97}$
60 0.065 2.1 0.0005 0.5 0.02 0.002 0.9992 $1.658\times 10^{-80}$
60 0.070 2.1 0.0005 0.5 0.02 0.002 1.0027 $3.675\times 10^{-77}$
60 0.075 2.1 0.0005 0.5 0.02 0.002 1.0049 $3.518\times 10^{-74}$
55 0.044 1.2 0.0002 2.5 0.02 0.002 0.9505 $9.349\times 10^{-100}$
55 0.044 1.2 0.0002 2.5 0.02 0.002 0.9586 $5.638\times 10^{-98}$
55 0.044 1.2 0.0002 2.5 0.02 0.002 0.9659 $3.541\times 10^{-96}$
55 0.065 1.2 0.0002 2.5 0.02 0.002 0.9989 $3.203\times 10^{-82}$
55 0.070 1.2 0.0002 2.5 0.02 0.002 1.0023 $6.647\times 10^{-79}$
55 0.075 1.2 0.0002 2.5 0.02 0.002 1.0045 $5.945\times 10^{-76}$
50 0.044 0.9 0.0001 2.5 0.02 0.002 0.9452 $9.367\times 10^{-99}$
50 0.046 0.9 0.0001 2.5 0.02 0.002 0.953 $4.892\times 10^{-97}$
50 0.048 0.9 0.0001 2.5 0.02 0.002 0.9602 $2.584\times 10^{-95}$
50 0.065 0.9 0.0001 2.5 0.02 0.002 0.9960 $6.235\times 10^{-82}$
50 0.070 0.9 0.0001 2.5 0.02 0.002 1.0003 $1.061\times 10^{-78}$
50 0.075 0.9 0.0001 2.5 0.02 0.002 1.0003 $8.316\times 10^{-76}$
45 0.044 0.9 0.00009 4.5 0.02 0.002 0.9587 $1.621\times 10^{-101}$
45 0.046 0.9 0.00009 4.5 0.02 0.002 0.99648 $2.331\times 10^{-99}$
45 0.048 0.9 0.00009 4.5 0.02 0.002 0.9703 $2.797\times 10^{-97}$
45 0.065 0.9 0.00009 4.5 0.02 0.002 0.9974 $9.544\times 10^{-83}$
45 0.070 0.9 0.00009 4.5 0.02 0.002 1.001 $1.876\times 10^{-79}$
45 0.075 0.9 0.00009 4.5 0.02 0.002 1.003 $1.571\times 10^{-76}$
Table 5: The table shows numerical values of $Q$ for different values of the parameters.
  $N$   $\Delta$   $C_{1}$   $\gamma$   $m$   $\alpha_{0}$   $\beta_{0}$   $Q$
65 0.048 2.2 0.0004 1 0.02 0.002 22.2046
65 0.050 2.2 0.0004 1 0.02 0.002 37.935
65 0.058 2.2 0.0004 1 0.02 0.002 282.574
60 0.048 2.2 0.0004 1 0.02 0.002 40.4028
60 0.050 2.2 0.0004 1 0.02 0.002 63.3381
60 0.058 2.2 0.0004 1 0.02 0.002 380.529
55 0.048 2.2 0.0004 1 0.02 0.002 90.3431
55 0.050 2.2 0.0004 1 0.02 0.002 127.607
55 0.058 2.2 0.0004 1 0.02 0.002 571.032
50 0.048 2.2 0.0004 1 0.02 0.002 259.167
50 0.050 2.2 0.0004 1 0.02 0.002 326.624
50 0.058 2.2 0.0004 1 0.02 0.002 1012.89
45 0.048 2.2 0.0004 1 0.02 0.002 976.335
45 0.050 2.2 0.0004 0.5 0.02 0.002 1101.49
45 0.058 2.2 0.0004 0.5 0.02 0.002 2277.39
Table 6: The table shows numerical results of $\frac{T}{H}$ for different values of the parameters.
  $N$   $\Delta$   $C_{1}$   $\gamma$     $\alpha_{0}$   $\beta_{0}$   $\frac{T}{H}$
65 0.082 2.2 0.0004 2.06 0.009 $4.560\times 10^{6}$
65 0.084 2.2 0.0004 2.06 0.009 $4.746\times 10^{6}$
65 0.086 2.2 0.0004 2.06 0.009 $4.916\times 10^{6}$
60 0.082 2.2 0.0004 2.06 0.009 $5.610\times 10^{6}$
60 0.084 2.2 0.0004 2.06 0.009 $5.870\times 10^{6}$
60 0.086 2.2 0.0004 2.06 0.009 $6.111\times 10^{6}$
55 0.082 2.2 0.0004 2.06 0.009 $6.901\times 10^{6}$
55 0.084 2.2 0.0004 2.06 0.009 $7.258\times 10^{6}$
55 0.086 2.2 0.0004 2.06 0.009 $7.593\times 10^{6}$
50 0.082 2.2 0.0004 2.06 0.009 $8.485\times 10^{6}$
55 0.084 2.2 0.0004 2.06 0.009 $8.97\times 10^{6}$
55 0.086 2.2 0.0004 2.06 0.009 $9.436\times 10^{6}$
45 0.082 2.2 0.0004 2.06 0.009 $1.041\times 10^{7}$
45 0.084 2.2 0.0004 2.06 0.009 $1.107\times 10^{7}$
45 0.086 2.2 0.0004 2.06 0.009 $1.171\times 10^{7}$
@@ 4 Discussion and Conclusion
Warm inflation offers a framework for comprehending the dynamics of the early universe in contrast to classical cold inflation. The inflationary paradigm is introduced to rectify the shortcomings of the traditional cosmological model. Inflationary scenarios can be classified into two categories: warm inflation and cold inflation. In cold inflation, the matter field does not interact with radiation and it slowly rolls to its flat potential. However, for the warm inflationary scenario, the inflaton interacts with other fields resulting in the transmission of energy from the inflaton to the radiation field during slow-roll. The inflaton completely decays into radiation when inflation comes to an end, preventing inflation from causing the universe to enter a very cold phase. As a result, the universe enters into radiation-dominated phase without the need for a separate reheating phase.
In this article, we have studied an inflationary scenario assuming that a holographic dark fluid is the source of inflation. We have chosen Barrow holographic dark energy for this purpose in a scenario where holographic dark energy interacts with radiation and energy transmits from holographic dark energy to radiation. Here we have considered the inflationary scenario in the high dissipative regime ($Q>1$). Assuming this condition we have reconstructed the Hubble parameter as a function of the e-folding number $N$. Slow roll conditions play an important role in warm inflation. A collection of slow-roll parameters determines how consistent the slow-roll approximation is. We have shown that these parameters will satisfy the warm inflationary conditions to validate our model. Moreover, after checking the tendencies of the different inflationary parameters, it was found that there is a good agreement with the observational data [79].
@@
In a concluding note we would like to do a short discussion regarding a study presented in Ref. [84] and why this is important for our study. In the Ref. [84] the authors discuss the possibility of a warm inflation scenario at a generalized level and show that it is very difficult and perhaps impossible to experience inflation driven by thermal effects. The paper shows that extreme fine-tuning is required between the energy released during the expansion of the universe and the number of particles created during the warm inflation to achieve that. The authors used a complete field theoretic approach and some rigorous methods of non-equilibrium quantum statistics to show that it is very difficult to realize such a scenario. In our work, we have followed a totally different avenue to study warm inflation. It should be noted here that the main stress in our work is the application of holography to realize warm inflation which was not the case in Ref. [84]. It is known that holography has its roots in black hole thermodynamics. So whenever such a concept is applied to any scenario, it is expected to bring out results that will involve stringent corrections. We think that in our work this has happened. The application of holography along with the unorthodox Barrow entropy (motivated by the fractal structure of the Covid-19 virus) may have introduced significant corrections to overhaul most of the difficulties discussed in Ref. [84] to realize warm inflation. This is the reason we have observationally favorable results (observationally favorable values of the inflationary parameters $\epsilon_{1}$, $r$, $n_{s}$, etc.). Moreover, there is enough support from the literature regarding the possibility of realizing a warm inflationary scenario in the presence of a holographic dark energy. In Ref. [74] the author has used the standard holographic dark energy to realize a warm inflationary scenario successfully. Other relevant works on warm inflation are [81, 85, 86]. These works provide enough support and motivation to explore a warm inflationary mechanism driven by holographic dark energy even when we have the knowledge presented in [84].
In future works, it would be interesting to examine our warm inflation model with Barrow holographic dark energy in the context of the Swampland conjectures. These conjectures, which provide criteria for effective field theories that can be consistently embedded in quantum gravity, have posed significant challenges for many inflationary models [87, 88]. Warm inflation has shown promise in addressing some of these challenges [89, 90, 91, 92, 93], particularly regarding the de Sitter conjecture, due to its distinct dynamics and reduced requirement for slow-roll. The fractal nature of Barrow entropy could provide additional interesting features when examined under Swampland constraints. Specifically, studying how the Barrow parameter $\Delta$ affects the satisfaction of various Swampland criteria, could offer new insights into the quantum gravitational consistency of warm inflationary models.
@@ Acknowledgments
P.R. acknowledges the Inter-University Centre for Astronomy and Astrophysics (IUCAA), Pune, India for granting visiting associateship. The authors acknowledge the hospitality and the research facilities of IUCAA during their scientific visits. The authors also thank the anonymous referee for his/her invaluable comments that helped them to improve the quality of the manuscript.
Data Availability Statement
No data was generated or analyzed in this study.
Conflict of Interest
There are no conflicts of interest.
Funding Statement
There is no funding to report for this article.